# Electric field tunable magnetoexcitons in Xenes-hBN-TMDC, Xenes-hBN-BP, and Xenes-hBN-TMTC heterostructures

### v1

@@ Abstract
In this work we propose novel van der Waals (vdW) heterostructures composed of Xenes, transition metal dichalcogenides (TMDCs), phosphorene, and transition metal trichalcogenides (TMTCs), which are separated by insulating hexagonal boron nitride (hBN) layers. We investigate theoretically the behavior of Rydberg indirect excitons in Xenes-hBN-TMDC, Xenes-hBN-BP, and Xenes-hBN-TMTC heterostructures, subject to parallel external electric and magnetic fields that are oriented perpendicular to the layers. By incorporating both isotropic and anisotropic materials, we demonstrate that excitonic properties can be effectively tuned through the external field strengths and the heterostructure design.
Our results show that the exciton reduced mass and the binding energy increase with the electric field strength, while enhanced dielectric screening from additional hBN layers reduces the binding energy. Anisotropic materials exhibit distinct excitonic responses, including variations in diamagnetic behavior. Moreover, the diamagnetic energy contributions and coefficients decrease with stronger electric fields but increase with the number of hBN layers. Finally, we explore the potential of time-periodic electric fields with Floquet band-structure engineering. These findings provide a comprehensive framework for controlling excitonic phenomena in low-dimensional materials, enabling the design of advanced optoelectronic and quantum devices.
@@ I Introduction
Two-dimensional layered materials have attracted persistent attention and have been extensively studied over the past two decades. Graphene has sparked significant research interest in two-dimensional (2D) layered materials such as transition metal dichalcogenides (TMDCs) [1], and group IVA materials such as silicene, germanene, and stanene, collectively known as Xenes, which consist of atoms arranged in a honeycomb lattice similar to that of graphene, but with varying degrees of buckling of two sublayers with respect to each other [2, 3]. In contrast to graphene, these materials exhibit either direct or indirect band gaps and possess remarkable physical properties as a result of their crystal symmetry and reduced dimensionality. Moreover, TMDCs and Xenes typically display in-plane structural isotropy in their electrical and optical properties.
@@
The concept of exciton formation in systems with spatially separated electrons and holes—specifically in coupled quantum wells where electrons and holes occupy different wells—was first introduced in Ref. [22], inspiring a wealth of theoretical and experimental studies on double quantum wells and heterostructures composed of two 2D materials. Today, heterostructures made from layers of TMDCs, phosphorene, Xenes, TMTCs, and others are at the forefront of nanomaterials research. The formation of excitonic complexes such as excitons, trions, and biexcitons in these systems leads to unique electronic, optical, and mechanical properties, which differ significantly from their bulk counterparts and hold great promise for applications in nanoelectronics, optoelectronics, and quantum computing [23, 24].
@@
Heterostructures comprising 2D monolayers separated by dielectric layers—particularly when there is a lattice mismatch between the constituent layers, as, for example, in MoS2–hBN–WSe2 or MoSe2–hBN–WSe2—have been extensively studied over the past two decades. These investigations have uncovered rich physics arising from interlayer coupling, moiré superlattices, and excitonic phenomena, providing deep insights into their electronic properties and excitonic behaviors, see [25, 26, 27] and references herein. Such studies underscore the critical role of lattice mismatch in designing and understanding the properties of 2D heterostructures, paving the way for novel applications in optoelectronics and quantum devices. Combining different 2D materials into heterostructures enables the tuning of electronic and optical properties via a precise control over stacking sequences, twist angles, and interlayer interactions. For example, a relative twist between layers can lead to the formation of moiré patterns that significantly influence the electronic band structure and may give rise to phenomena such as superconductivity [28] and flat electronic bands [29].
@@
The band structures of TMDC, phosphorene, and TMTC monolayers are largely unaffected by a perpendicular electric field, and the same holds true for heterostructures composed of these monolayers. In contrast, Xenes exhibit a tunable band gap due to their buckled structure. Therefore, heterostructures such as Xenes-hBN-TMDC, Xenes-hBN-BP, or Xenes-hBN-TMTC can exhibit a tunable gap, which in turn affects the properties of indirect excitons formed within them. The buckled geometry of Xenes leads to a potential difference between sublattices, allowing control over the band gap, effective masses of charge carriers, binding energies, and diamagnetic coefficients (DMCs) of magnetoexcitons. We do not consider systems where Xene monolayers are placed directly on top other layers without hBN separation, as Xenes are unstable in air [30, 31].
@@
In 2D systems, the reduced dielectric screening and strong Coulomb interactions enhance the binding of excitons, making them especially sensitive to magnetic fields. When a magnetic field is applied, it affects both the binding energy and the diamagnetic shift of the exciton. In van der Waals heterostructures, composed of novel 2D materials, the impact of a magnetic field on excitons is further influenced by the nature of interlayer coupling, effective masses, dielectric environments. These systems allow for tunable magnetoexcitonic behavior through effective mass anisotropy [32, 8], layer engineering, electric field control.
@@
The goals of this article are twofold: (i) to demonstrate the tunability of the properties of indirect excitons in Rydberg states within isotropic (Xenes-hBN-TMDC) and anisotropic (Xenes-hBN-BP or Xenes-hBN-TMTC) heterostructures using external electric and magnetic fields; and (ii) to explore the use of a time-periodic electric field as a tool for band-structure engineering. The remainder of this article is organized as follows. In Section II, we present the theoretical model used to describe the electron-hole system in the presence of external electric and magnetic fields, both with and without effective mass anisotropy of the charge carriers. Section III contains the results of our calculations for the binding energies (BEs) of Rydberg exciton states under an external electric field in vdW heterostructures. In the same section, we also present and discuss the BEs and diamagnetic coefficients (DMCs) of magnetoexcitons in these heterostructures under the influence of an external electric field. Additionally, we explore a band-structure engineering via a time-periodic electric field. Finally, our concluding remarks are given in Section IV.
Figure 1: (Color online) Schematic illustration of excitons in ($a$) Xenes-hBN-TMDC and ($b$) Xenes-hBN-BP van der Waals heterostructure in electric and magnetic fields. By replacing phosphorene by TMTC monolayer one gets Xenes-hBN-TMTC heterostructure.
@@ II Theoretical Model
It is known that electrostatically bound electrons and holes in an external magnetic field form magnetoexcitons. In this section, following Refs. [33, 34, 35, 36], we briefly introduce the theoretical model for describing Mott-Wannier magnetoexcitons in Xenes-hBN-TMDC and Xenes-hBN-BP or Xenes-hBN-TMTC heterostructures. We consider the energy contribution, $\Delta E$, from external electric and magnetic fields to the binding energies, $E_{B}$, of Rydberg states of magnetoexcitons and their diamagnetic coefficients (DMCs). In the system under consideration, excitons are confined within a van der Waals heterostructure, where $N$ layers of hBN monolayers separate Xenes and TMDC or Xenes and phosphorene or Xenes and TMTC monolayers, respectively.
Let us outline the low-energy model that describes exciton states in Xenes-hBN-TMDC and Xenes-hBN-BP van der Waals heterostructures in the $x-y$ plane. We consider an electron and a hole in parallel magnetic $\mathbf{B}=(0,0,B_{z})\equiv(0,0,B)$ and electric $\mathbf{E}=(0,0,E_{z})\equiv(0,0,E)$ fields, which are perpendicular to the heterostructure, as depicted in Fig. 1. Monolayers of silicene, germanene, and low-buckled stanene can be pictured as honeycomb graphene monolayers with an out-of-plane buckling, such that the A and B triangular sublattices are offset with respect to the $x-y$ plane with a typical distance $d_{0}$ to the latter. This distance is known as the buckling constant or buckling factor. The intrinsic sensitivity of Xenes to a perpendicular electric field is due to the offset between the two triangular sublattices. In particular, the resulting asymmetry causes an on-site potential difference between the sublattices. In the absence of an external electric field the band structure of Xenes in the vicinity of the $K/K^{\prime}$ points resembles gapped graphene, while the application of a perpendicular electric field changes the band gap in the monolayer Xenes. hus, the single-particle dispersion of quasiparticles in a monolayer Xenes is described in the vicinity of the $K/K^{\prime}$ points by the two-dimensional (2D) Hamiltonian [37]
@@
where ${\hat{\sigma}}_{j}$ and ${\hat{\tau}}_{j}$ are Pauli matrices in the spin and in the pseudospin space, respectively. The pseudospin is a result of the underlying honeycomb lattice. $\xi=\pm 1$ refers to the valley index of the honeycomb lattice. Moreover, $v_{F}$ is the Fermi velocity, $\Delta_{so}$ is the spin-orbit splitting, and $\Delta_{z}=ed_{0}E$ describes the potential caused by a perpendicular electric field $E$ on the two triangular sublattices of the honeycomb lattice. Since only ${\hat{\sigma}}_{z}$ appears, the Hamiltonian $H_{0}$ separates in the spin space into two massive 2D Dirac Hamiltonian with dispersions
@@
with $\Delta_{\xi\sigma}=|\xi\sigma\Delta_{so}-ed_{0}E|$, where $\xi,\sigma=\pm 1$ are the valley and spin indices, respectively. Thus, we have four bands with two different gaps $\Delta_{\xi\sigma}$ due to $\xi\sigma=\pm 1$. The shift of $\Delta_{\xi\sigma}$ by the electric field $E$ was previously discussed in Refs. [38, 39, 40, 41], as well as in Ref. [37]. For the following we assume that the kinetic energy $v_{F}^{2}p^{2}$ is small compared to $\Delta_{\xi\sigma}^{2}$, such that we can expand the square root to obtain a parabolic dispersion $E_{\sigma}(p)\sim\pm(|\Delta_{\xi\sigma}|+p^{2}/2m)$ with the effective electron mass $m=|\Delta_{\xi\sigma}|/v_{F}^{2}$ [42, 2]. Thus, the conduction and valence bands are parabolic in the vicinity of the $K/K^{\prime}$ points. The effective masses of electrons and holes are the same due to the symmetry between the lowest conduction and highest valence bands.
At non-zero electric fields, both the valence and conduction bands, split into upper bands with a large gap (when $\xi=-\sigma$), and lower bands with a small gap (when $\xi=\sigma$). We call the excitons formed by charge carriers from the large gap $A$ excitons, and those formed by charge carriers in the small gap $B$ excitons. At small electric fields, germanene and especially stanene show significant differences between the masses of the $A$ and $B$ excitons. The latter is due to their large intrinsic band gaps. Silicene, which has an intrinsic band gap on the order of a couple of meV, exhibits very little difference between the masses of $A$ and $B$ excitons, even at relatively small electric fields. At large electric field the difference between the $A$ and $B$ exciton mass is negligible in silicene and germanene. In all cases, the mass of the $A$ exciton exceeds the mass of the $B$ exciton. TMDC monolayers are isotropic and effective masses of an electron and hole are different and $m_{h}>m_{e}$.
In the next step we will consider a system whose translation invariance is broken by an external magnetic field. To this end, let us introduce the coordinate vectors of the electron and hole for the Mott-Wannier exciton in the Xenes and TMDC (phosphorene or TMTC) layers. The following in-plane coordinates $\mathbf{r}_{e}(x_{e},y_{e})$ for an electron in Xenes layer and $\mathbf{r}_{h}(x_{h},y_{h})$ for a hole in TMDC or phosphorene layer, respectively, are used in our description. The electron/hole spins are polarized by the magnetic field and the honeycomb structure is irrelevant because the system is dominated by the magnetic length. Then the Hamiltonian of an interacting pair of an electron at the site $r_{e}$ and a hole at the site $r_{h}$ reads ($\hbar=c=1$)
@@
where $V\left(|\mathbf{r}_{e}-\mathbf{r_{h}}|\right)$ describes the Coulomb interaction between the electron and hole via a cental Coulomb potential
@@
where $r^{2}=x^{2}+y^{2}$. The effective masses $m^{e}$ and $m^{h}$ are taken from the calculation of the expanded square-root dispersion of Eq. (2). The anisotropic nature of the phosphorene and TMTC, in contrast to other 2D isotropic materials such as graphene and TMDC semiconductors, breaks the circular symmetry. This requires the use of Cartesian coordinates for the description of excitons. The asymmetry of the electron and hole dispersion in phosphorene is reflected in the Hamiltonian for the Mott-Wannier exciton, and the Hamiltonian for an interacting electron-hole pair in Xenes-hBN-BP or Xenes-hBN-TMTC heterostructures in the external electric and magnetic fields is obtained within the framework of the effective mass approximation. In particular, the electric field dependence of Hamiltonian (3) is implicit via the effective mass of electron. Equation  (4) describes the interaction between the electron and the hole that are located in different Xenes and phosphorene or TMTC monolayers, separated by a distance $D=h+Nl_{\text{hBN}}$. $l_{\text{hBN}}=0.333$ nm is the thickness of the hBN layer and $h$ is the thickness of the phosphorene or TMTC, while $m_{x}^{h}$ and $m_{y}^{h}$ correspond to the effective mass of the hole in the $x$ or $y$ direction in the phosphorene or TMTC, respectively. In the case of Xenes TMDC heterostructure $m_{x}^{e}=m_{x}^{h}=m_{e}$ and the Hamiltonian (3) becomes simpler.
Following the standard procedure [43] for the separation of the relative motion of the electron-hole pair from their center-of-mass motion, one introduces variables for the center-of-mass of an electron-hole pair. After a lengthy calculation, following Refs. [44, 34, 35, 36], one obtains the equation that describes the Mott–Wannier exciton in the Rydberg optical states in the external electric and magnetic field perpendicular to the Xenes-hBN-BP heterostructure. Finally, the equation for the relative motion of the electron and hole in the Xenes-hBN-BP heterostructure with zero center-of-mass momentum [33, 44, 45] reads
@@
where $\mu_{x}=\frac{m^{e}m_{x}^{h}}{m^{e}+m_{x}^{h}}$ and $\mu_{y}=\frac{m^{e}m_{y}^{h}}{m^{e}+m_{y}^{h}}$ are the reduced masses in the $x$ and $y$ directions, respectively. In Eq. (5) the anisotropy is present in the kinetic and magnetic terms, while the potential term has isotropic form and the action of the electric field is present via anisotropic reduced masses $\mu_{x}$ and $\mu_{y}$. We solve Eq. (5) using $V\left(\sqrt{x^{2}+y^{2}+D^{2}}\right)$ for indirect magnetoexcitons. For indirect magnetoexcitons, both $V_{RK}$ and $V_{C}$ potentials are used to investigate the importance of using correct potential describing electron-hole interactions. Note that Eq. (5) does not explicitly contain any spin- or valley-dependent Zeeman terms.
The Schrödinger equation with Hamiltonian (3) for Xenes-hBN-TMDC heterostructure when $m_{x}^{e}=m_{x}^{h}=m_{h}$ has the form: $\hat{H}\Psi(\mathbf{r}_{e},\mathbf{r}_{h})=E\Psi(\mathbf{r}_{e},\mathbf{r}_{h})$, where $\Psi(\mathbf{r}_{e},\mathbf{r}_{h})$ and $E$ are the eigenfunction and eigenenergy, respectively. Due to the spherical symmetry, in this case the corresponding equation easier write in spherical coordinates. Finally, after separating the angular variable the equation with zero center-of-mass momentum reads [44, 45]:
$\left[-\frac{1}{2\mu}\frac{{\partial}^{2}}{\partial r^{2}}-\frac{1}{2\mu}\frac{1}{r}\frac{{\partial}}{\partial r}+\frac{e^{2}}{8\mu}B^{2}r^{2}+V(r)\right]\Phi(r)=E\Phi(r).$ (6)
where $\mu=\frac{m_{e}m_{h}}{m_{e}+m_{h}}$ is the reduced mass. Equation (5) describes the Mott–Wannier magnetoexciton in Rydberg optical states in 2D materials. This equation has a long history of the solution in the case of the electron-hole Coulomb interaction [46, 33, 47, 48, 49, 50, 51, 44, 52, 45].

### v2

# Electric field tunable magnetoexcitons in Xenes/hBN/TMDC, Xenes/hBN/BP, and Xenes/hBN/TMTC heterostructures

@@ Abstract
In this work, we propose novel van der Waals (vdW) heterostructures composed of Xenes, transition metal dichalcogenides (TMDCs), phosphorene, and transition metal trichalcogenides (TMTCs), which are separated by insulating hexagonal boron nitride (hBN) layers. We theoretically investigate the behavior of Rydberg indirect excitons in Xenes/hBN/TMDC, Xenes/hBN/BP, and Xenes/hBN/TMTC heterostructures subjected to parallel external electric and magnetic fields that are oriented perpendicular to the layers. By incorporating both isotropic and anisotropic materials, we demonstrate that excitonic properties can be effectively tuned through the external field strengths and the heterostructure design.
Our results show that the exciton reduced mass and the binding energy increase with the electric field strength, while enhanced dielectric screening from additional hBN layers reduces the binding energy. Anisotropic materials exhibit distinct excitonic responses, including variations in diamagnetic behavior. Moreover, the diamagnetic energy contributions and coefficients decrease with stronger electric fields but increase with the number of hBN layers. Finally, we explore the potential of time-periodic electric fields with Floquet band-structure engineering. These findings provide a comprehensive framework for controlling excitonic phenomena in low-dimensional materials.
@@ I Introduction
Two-dimensional layered materials have attracted persistent attention and have been extensively studied over the past two decades. Graphene has sparked significant research interest in two-dimensional (2D) layered materials such as transition metal dichalcogenides (TMDCs) [1] and group IVA materials such as silicene, germanene, and stanene, collectively known as Xenes, which consist of atoms arranged in a honeycomb lattice similar to that of graphene, but with varying degrees of buckling of two sublayers with respect to each other [2, 3]. In contrast to graphene, these materials exhibit either direct or indirect band gaps and possess remarkable physical properties as a result of their crystal symmetry and reduced dimensionality. Moreover, TMDCs and Xenes typically display in-plane structural isotropy in their electrical and optical properties.
@@
The concept of exciton formation in systems with spatially separated electrons and holes—specifically in coupled quantum wells where electrons and holes occupy different wells—was first introduced in Ref. [22], inspiring a wealth of theoretical and experimental studies on double quantum wells and heterostructures composed of 2D materials. Today, heterostructures made from layers of TMDCs, phosphorene, Xenes, TMTCs, and others are at the forefront of nanomaterials research. The formation of excitonic complexes such as excitons, trions, and biexcitons in these systems leads to unique electronic, optical, and mechanical properties, which differ significantly from their bulk counterparts and hold great promise for applications in nanoelectronics, optoelectronics, and quantum computing [23, 24].
@@
Heterostructures comprising 2D monolayers separated by dielectric layers—particularly when there is a lattice mismatch between the constituent layers, as, for example, in MoS2/hBN/WSe2 or MoSe2/hBN/WSe2—have been extensively studied over the past two decades. These investigations have uncovered rich physics arising from interlayer coupling, moiré superlattices, and excitonic phenomena, providing deep insights into their electronic properties and excitonic behaviors, see [25, 17, 26] and references herein. Combining different 2D materials into heterostructures enables the tuning of electronic and optical properties [27, 28].
@@
The band structures of TMDC, phosphorene, and TMTC monolayers are largely unaffected by a perpendicular electric field, and the same holds true for heterostructures composed of these monolayers. In contrast, Xenes exhibit a tunable band gap due to their buckled structure. Therefore, heterostructures such as Xenes/hBN/TMDC, Xenes/hBN/BP, or Xenes/hBN/TMTC can exhibit a tunable gap, which in turn affects the properties of indirect excitons formed within them. The buckled geometry of Xenes leads to a potential difference between sublattices, allowing control over the band gap, effective masses of charge carriers, binding energies (BEs), and diamagnetic coefficients (DMCs) of magnetoexcitons. We do not consider systems where Xene monolayers are placed directly on top of other layers without hBN separation, as Xenes are unstable in air [29, 30].
@@
In 2D systems, the reduced dielectric screening and strong Coulomb interactions enhance the binding of excitons, making them especially sensitive to magnetic fields. When a magnetic field is applied, it affects both the binding energy and the diamagnetic shift of the exciton. In van der Waals heterostructures, composed of novel 2D materials, the impact of a magnetic field on excitons is further influenced by the nature of interlayer coupling, effective masses, and dielectric environments. These systems allow for tunable magnetoexcitonic behavior through effective mass anisotropy [31, 8], layer engineering, and electric field control.
@@
The goals of this article are twofold: (i) to demonstrate the tunability of the properties of indirect excitons in Rydberg states within isotropic (Xenes/hBN/TMDC) and anisotropic (Xenes/hBN/BP or Xenes/hBN/TMTC) heterostructures using external electric and magnetic fields; and (ii) to explore the use of a time-periodic electric field as a tool for band-structure engineering. The remainder of this article is organized as follows. In Section II, we present the theoretical model used to describe the electron-hole system in the presence of external electric and magnetic fields, both with and without effective mass anisotropy of the charge carriers. Section III contains the results of our calculations for the binding energies of Rydberg exciton states under an external electric field in vdW heterostructures. In the same section, we also present and discuss the BEs and DMCs of magnetoexcitons in these heterostructures under the influence of an external electric field. Additionally, we explore a band-structure engineering via a time-periodic electric field. Finally, our concluding remarks are given in Section IV.
Figure 1: (Color online) Schematic illustration of excitons in ($a$) Xenes/hBN/TMDC and ($b$) Xenes/hBN/BP van der Waals heterostructure in electric and magnetic fields. By replacing phosphorene by TMTC monolayer one gets Xenes/hBN/TMTC heterostructure.
@@ II Theoretical Model
It is known that electrostatically bound electrons and holes in an external magnetic field form magnetoexcitons. In this section, following Refs. [32, 33, 34, 35], we briefly introduce the theoretical model for describing Mott-Wannier magnetoexcitons in Xenes/hBN/TMDC and Xenes/hBN/BP or Xenes/hBN/TMTC heterostructures. We consider the energy contribution, $\Delta E$, from external electric and magnetic fields to the binding energies, $E(B)$, of Rydberg states of magnetoexcitons and their diamagnetic coefficients. In the system under consideration, excitons are confined within a van der Waals heterostructure, where $N$ layers of hBN monolayers separate Xenes and TMDC or Xenes and phosphorene or Xenes and TMTC monolayers.
Let us outline the low-energy model that describes exciton states in Xenes/hBN/TMDC and Xenes/hBN/BP van der Waals heterostructures in the xyx-y plane. We consider an electron and a hole in parallel magnetic $\mathbf{B}=(0,0,B_{z})\equiv(0,0,B)$ and electric $\mathbf{E}=(0,0,E_{z})\equiv(0,0,E)$ fields, which are perpendicular to the heterostructure, as depicted in Fig. 1. Monolayers of silicene, germanene, and low-buckled stanene can be pictured as honeycomb graphene monolayers with an out-of-plane buckling, such that the A and B triangular sublattices are offset with respect to the xyx-y plane with a typical distance $d_{0}$ to the latter. This distance is known as the buckling constant or buckling factor. The intrinsic sensitivity of Xenes to a perpendicular electric field is due to the offset between the two triangular sublattices. In particular, the resulting asymmetry causes an on-site potential difference between the sublattices. In the absence of an external electric field the band structure of Xenes in the vicinity of the $K/K^{\prime}$ points resembles gapped graphene, while the application of a perpendicular electric field changes the band gap in the monolayer Xenes.Thus, the single-particle dispersion of quasiparticles in a monolayer Xenes is described in the vicinity of the $K/K^{\prime}$ points by the two-dimensional (2D) Hamiltonian [36]
@@
where ${\hat{\sigma}}_{j}$ and ${\hat{\tau}}_{j}$ are Pauli matrices in the spin and in the pseudospin space, respectively. The pseudospin is a result of the underlying honeycomb lattice. $\xi=\pm 1$ refers to the valley index of the honeycomb lattice. Moreover, $v_{F}$ is the Fermi velocity, $\Delta_{so}$ is the spin-orbit splitting, and $\Delta_{z}=ed_{0}E_{\perp}$ describes the potential caused by a perpendicular electric field $E_{\perp}$ on the two triangular sublattices of the honeycomb lattice. Since only ${\hat{\sigma}}_{z}$ appears, the Hamiltonian $H_{0}$ separates in the spin space into two massive 2D Dirac Hamiltonian with dispersions
@@
with $\Delta_{\xi\sigma}=|\xi\sigma\Delta_{so}-ed_{0}E_{\perp}|$, where $\xi,\sigma=\pm 1$ are the valley and spin indices, respectively. Thus, we have four bands with two different gaps $\Delta_{\xi\sigma}$ due to $\xi,\sigma=\pm 1$. The shift of $\Delta_{\xi\sigma}$ by the electric field $E_{\perp}$ was previously discussed in Refs. [37, 38, 39, 40], as well as in Ref. [36]. For the following, we assume that the kinetic energy $v_{F}^{2}p^{2}$ is small compared to $\Delta_{\xi\sigma}^{2}$, such that we can expand the square root to obtain a parabolic dispersion $E_{\sigma}(p)\sim\pm(|\Delta_{\xi\sigma}|+p^{2}/2m)$ with the effective electron mass $m=|\Delta_{\xi\sigma}|/v_{F}^{2}$ [41, 2]. Thus, the conduction and valence bands are parabolic in the vicinity of the $K/K^{\prime}$ points. The effective masses of electrons and holes are the same due to the symmetry between the lowest conduction and highest valence bands, and can be written as a function of the external electric field in the following form:
$m^{e}=\frac{\lvert\xi\sigma\Delta_{so}-ed_{0}E_{\perp}\rvert}{v_{F}^{2}}.$ (3)
At non-zero electric fields, both the valence and conduction bands, split into upper bands with a large gap (when $\xi=-\sigma$), and lower bands with a small gap (when $\xi=\sigma$). We call the excitons formed by charge carriers from the large gap $A$ excitons, and those formed by charge carriers in the small gap $B$ excitons. At small electric fields, germanene and especially stanene show significant differences between the masses of the $A$ and $B$ excitons. The latter is due to their large intrinsic band gaps. Silicene, which has an intrinsic band gap on the order of a couple of meV, exhibits very little difference between the masses of $A$ and $B$ excitons, even at relatively small electric fields. At large electric field the difference between the $A$ and $B$ exciton mass is negligible in silicene and germanene. In all cases, the mass of the $A$ exciton exceeds the mass of the $B$ exciton. TMDC monolayers are isotropic and effective masses of an electron and hole are different.
In the next step, we will consider a system whose translation invariance is broken by an external magnetic field. To this end, let us introduce the coordinate vectors of the electron and hole for the Mott-Wannier exciton in the Xenes and TMDC (phosphorene or TMTC) layers. The following in-plane coordinates $\mathbf{r}_{e}(x_{e},y_{e})$ for an electron in Xenes layer and $\mathbf{r}_{h}(x_{h},y_{h})$ for a hole in TMDC or phosphorene layer, respectively, are used in our description. The electron/hole spins are polarized by the magnetic field, and the honeycomb structure is irrelevant because the system is dominated by the magnetic length. Then the Hamiltonian of an interacting pair of an electron at the site $r_{e}$ and a hole at the site $r_{h}$ reads ($\hbar=c=1$)
@@
where $m_{x}^{h}$ and $m_{y}^{h}$ correspond to the effective mass of the hole in the $x$ or $y$ direction in the phosphorene or TMTC, respectively.
In Eq. (4) $V\left(|\mathbf{r}_{e}-\mathbf{r_{h}}|\right)$ describes the Coulomb interaction between the electron and hole via a central Coulomb potential
@@
where $r^{2}=x^{2}+y^{2}$. The anisotropic nature of the phosphorene and TMTC, in contrast to other 2D isotropic materials such as graphene and TMDC semiconductors, breaks the circular symmetry. This requires the use of Cartesian coordinates for the description of excitons. The asymmetry of the electron and hole dispersion in phosphorene is reflected in the Hamiltonian for the Mott-Wannier exciton. The Hamiltonian for an interacting electron-hole pair in Xenes/hBN/BP or Xenes/hBN/TMTC heterostructures in the external electric and magnetic fields is obtained within the framework of the effective mass approximation. In particular, the electric field dependence of Hamiltonian (4) is implicit via the effective mass of electron in Xenes monolayer. Equation  (5) describes the interaction between the electron and the hole that are located in different Xenes and phosphorene or TMTC monolayers, separated by a distance $D=h+Nl_{\text{hBN}}$. $l_{\text{hBN}}=0.333$ nm is the thickness of the hBN layer, $h$ is the average thickness of the Xene and phosphorene or TMTC, and $N$ is the number of hBN layers.
Following the standard procedure [42] for the separation of the relative motion of the electron-hole pair from their center-of-mass motion, one introduces variables for the center-of-mass of an electron-hole pair. After a lengthy calculation, following Refs. [43, 33, 34, 35], one obtains the equation that describes the Mott–Wannier exciton in the Rydberg optical states in the external electric and magnetic field perpendicular to the Xenes/hBN/BP heterostructure. Finally, the equation for the relative motion of the electron and hole in the Xenes/hBN/BP heterostructure in the center-of-mass system (cf. Appendix and Refs. [32, 43, 44]) reads
@@
where
$\mu_{x}=\frac{m^{e}m_{x}^{h}}{m^{e}+m_{x}^{h}}\ \ \text{and}\ \ \mu_{y}=\frac{m^{e}m_{y}^{h}}{m^{e}+m_{y}^{h}}$ (7)
are the reduced masses in the $x$ and $y$ directions, respectively. In Eq. (6) the anisotropy is present in the kinetic and magnetic terms, while the potential term has isotropic form and the action of the electric field is present via anisotropic reduced masses $\mu_{x}$ and $\mu_{y}$. We solve Eq. (6) using $V\left(\sqrt{x^{2}+y^{2}+D^{2}}\right)$ for indirect magnetoexcitons. Note that Eq. (6) does not explicitly contain any spin- or valley-dependent Zeeman terms.
As discussed in the Appendix, the center-of-mass dynamics is described by a harmonic oscillator of mass $M$ and with frequency $\varpi=\frac{eB}{2\sqrt{M\mu}}$ for the isotropic system with $\mu_{x}=\mu_{y}$. Thus, the energy of the interacting electron–hole pair in a magnetic field is the sum of the energies of the relative motion and the center-of-mass motion. Notice that when the exciton is considered beyond the center-of-mass frame, the separation of collective (center-of-mass) and internal motions for two particles in a magnetic field becomes a long-standing and challenging problem  [45, 32, 46, 47, 48, 52, 49, 50, 43, 51, 44]. Such separation is only possible for two particles with equal masses and equal magnitudes of charge. In this special case, the center-of-mass motion becomes quantized and forms the Landau levels.
For excitons in our heterostructures, this separation is not possible. The Landau orbit of the center-of-mass and the internal excitonic motion are strongly coupled [53]. In Ref. [53], the authors investigate several physical scenarios in which this coupling becomes significant even at laboratory magnetic-field strengths. Notably, the Landau orbit itself can be substantially modified by this interaction. In contrast, for Rydberg exciton states, our system does not fall into that category. Therefore, for describing the internal exciton motion, we follow the approaches developed in [32, 43, 44].
The Schrödinger equation with Hamiltonian (4) for Xenes/hBN/TMDC heterostructure when for TMDC monolayer $m_{x}^{e}=m_{x}^{h}\equiv m^{h}$ has the form: $\hat{H}\Psi(\mathbf{r}_{e},\mathbf{r}_{h})=E\Psi(\mathbf{r}_{e},\mathbf{r}_{h})$, where $\Psi(\mathbf{r}_{e},\mathbf{r}_{h})$ and $E$ are the eigenfunction and eigenenergy, respectively. Due to the spherical symmetry, in this case, the corresponding equation easier write in spherical coordinates. Finally, after separating the angular variable the equation in the center-of-mass momentum system [43, 44]:
$\left[-\frac{1}{2\mu}\frac{{\partial}^{2}}{\partial r^{2}}-\frac{1}{2\mu}\frac{1}{r}\frac{{\partial}}{\partial r}+\frac{e^{2}}{8\mu}B^{2}r^{2}+V(r)\right]\Phi(r)=E\Phi(r),$ (8)
where
$\mu=\frac{m^{e}m^{h}}{m^{e}+m^{h}}$ (9)
is the reduced mass. This equation describes the Mott–Wannier magnetoexciton in Rydberg optical states in 2D materials. Here we have neglected the terms that couple the center-of-mass and the relative coordinates, an approximation that is validated within our calculation. Eq. (6) has a long history in the context of the electron-hole Coulomb interaction [45, 32, 46, 47, 48, 49, 50, 43, 51, 44].